If the the function is C(g)= 3.03 (g)
What is C(2)?
What is C(a)?
How do you solve this?

Answer by bucky(2189)

All these two problems are telling you to do is to replace the variable in the function with the quantity in the parentheses.
.
So for C(2) just replace g with 2.
.
Start with C(g) = 3.03(g)
.
For the first problem replace every g with 2. When you do that you get the equation:
.
C(2) = 3.03(2)
.
Multiply out the right side by multiplying 3.03 times 2 and the result is:
.
C(2) = 6.06 .... that's the first answer.
.
For the second problem, follow the same process, but replace every g with "a"
.
Again start with C(g) = 3.03(g)
.
Replace every g with a to get:
.
C(a) = 3.03(a)
.
And that's the answer to the second problem you were given.
.
Pretty easy once you understand what the problem is asking you to do.
.
Hope this helps.